OpenStudy (eric_d):

If A is an acute angle with cos A= 5/7 & B is an obtuse angle with sin B=1/5, find values of sin (A-B) and cos(A-B)

OpenStudy (dumbcow):

\[\sin(a-b) = \sin a \cos b - \sin b \cos a\] \[\cos(a-b) = \cos a \cos b + \sin a \sin b\]

OpenStudy (eric_d):

ok.. what shld I do next

OpenStudy (eric_d):

@dumbcow

OpenStudy (dumbcow):



OpenStudy (eric_d):